# Robust Superradiance and Spontaneous Spin Ordering in Disordered Waveguide QED

### v1

@@ VII Conclusions
In summary, we investigated the fate of Dicke superradiance in atomic arrays subject to strong disorder. Using semiclassical methods tailored for collective emission, we numerically demonstrate that the characteristic $N^{2}$ scaling of the peak emission rate and the scaling $t_{\star}\sim\log(N)/N$ of the peak emission time remain remarkably robust, even when atomic positions are strongly disordered. In the extreme limit of a homogeneously distributed atomic ensemble, we found that the peak rate is only reduced by a constant factor of approximately $1/4$ and that there is no change in the asymptotic scaling behavior. However, disorder affects the finite-size scaling toward this limit, which explains the difficulty of exploring these effects with most of the previously employed numerical methods that are limited to small atom numbers.
@@
To gain physical understanding of our numerical findings on this robustness, we developed a variational product-state ansatz that yields a physically motivated upper bound on the collective decay rate. The bound $\mathbb{R}_{<}$ in Eq. (33) not only captures the observed superradiant scaling for disorder of varying strength, but also offers conceptual insight into the mechanism behind its robustness. Specifically, it reveals a form of disorder-induced self-organization: instead of synchronizing to align along a global direction as in the weakly disordered case, the atomic spins align coherently along locally determined directions set by the relative propagation phases. This form of spontaneous spin ordering always leads to constructive interference for collective emission, even under strong disorders. Then, using QSDMF simulations that can provide access to trajectory-level quantities, the spin ordering is confirmed numerically. Interestingly, this spin ordering also reveals the mechanism that underlies the asymmetric photon correlations reported in [28], since the two ordering patterns prefer emission to the left and the right, respectively.
@@
Looking forward, several promising directions emerge. While our current model involves two decay channels arising from the bidirectional waveguide, extensions to systems with multiple decay channels may reveal new forms of collective phenomena. Additionally, studying superradiance in engineered environments—such as squeezed reservoirs, structured bath, or chiral waveguides—may uncover richer phenomena of coherence and self-organization. The analytical and numerical frameworks introduced here are broadly adaptable for future studies of disorder, coherence, and collectivity in many-body quantum optics.
Note added
Upon completing the writing of our manuscript, we became aware of a related work [70], which discusses spin synchronization in the steady states of spins coupled to ring cavities or waveguides.
@@ Acknowledgements.
Appendix A Two-Mode Structure and Finite-Size Scaling
In Sec. IV.1, we show a disorder-dependent finite-size scaling relation (18). To understand this behavior, we now examine the structure of the collective decay rate operator $\hat{\mathcal{R}}$ and derive a bound on the maximal decay rate. This bound reveals a two-mode structure in the decay channels and can give an understanding of the scaling crossover observed numerically.
For the collective decay rate $\mathcal{R}$, we know we can write
@@ Appendix A Two-Mode Structure and Finite-Size Scaling
where $\hat{R}$ is a Hermitian matrix defined by the jump operators in Eq. (3), $\Gamma_{k}$ are its eigenvalues, and $\Omega_{k}=\sum_{j}U_{kj}\sigma_{j}^{-}$ are the corresponding decay modes after diagonalization of $\hat{R}$ with a unitary transformation $U$. Let $\Gamma_{\pm}$ denote the two eigenvalues in the current case. Then, the maximal possible total decay rate, without any knowledge of the underlying dynamics, can be bounded as (see Appendix F)
>, $\mathcal{R}\leq\mathbb{R}_{>},\text{ with }\mathbb{R}_{>}=\frac{3}{2}N\Gamma_{\max}.$ (42)
Note that our product state ansatz works better for superradiance due to the consideration of the separability in superradiant dynamics. This thus gives a more loose upper bound on the collective decay rate, but it turns out to provide insight into the finite-size scaling.
@@
of which the diagonalization gives two eigenvalues $\Gamma_{\pm}=\frac{\gamma}{2}N(1\pm|c|)$. Then our original master equation (1) can be rewritten as (5) with two orthogonal modes.
Since $\xi_{j}$ is uniformly distributed in $\Theta/2,\Theta/2)$, we know $\Re e^{-i2\xi}$ and $\Im e^{-i2\xi}$ have the means and variances shown in Eq. (19). Thus, for large $N$, $c$ obeys a Gaussian distribution with $\Re c\sim N(\mu_{a},\sigma_{a}^{2}/N)$, $\Im c\sim N(\mu_{b},\sigma_{b}^{2}/N)$. In the thermodynamic limit $N\rightarrow\infty$, we get Eq. (20), which gives $\mathbb{R}_{>}\sim N^{2}$ for all disorder strength.
To probe the finite-size scaling, we consider large but finite $N$. To get an analytical understanding, we can write $\Re c=\mu_{a}+\delta_{a}$ and $\Im c=\delta_{b}$, with $\delta_{a/b}$ being small Gaussian fluctuations $\sim N^{-1/2}$. Then we see
@@
For $\Theta=n\pi$, since $\mu_{a}=0$, $|c|=\sqrt{(\delta_{a}^{2}+\delta_{b}^{2})}\sim|\delta_{a/b}|\sim N^{-1/2}$. That is, $E[\Gamma_{\max}]=\frac{1}{2}\gamma N\left(1+\mathcal{O}\left(N^{-1/2}\right)\right)$, which indeed agrees with the exact result (21). For $\Theta\neq n\pi$, since $\mu_{a}\neq 0$, we have the expansion
$|c|=\mu_{a}\left(1+\frac{\delta_{a}}{\mu_{a}}+\frac{1}{2}\frac{\delta_{b}^{2}}{\mu_{a}^{2}}+\ldots\right).$ (47)
We thus know
$E[|c|]=\mu_{a}+\frac{1}{2\mu_{a}}E[\delta_{b}^{2}]\sim\mu_{a}+\mathcal{O}(N^{-1}),$ (48)
which gives the scaling
$E[\Gamma_{\max}]=\frac{1}{2}\gamma N\left(1+\mu_{a}+\mathcal{O}\left(N^{-1}\right)\right).$ (49)
The finite-size correction to $\mathbb{R}_{>}/N^{2}$ is then of order $\mathcal{O}(N^{-1})$, which also agrees with our numerical observation. Therefore, the convergence of $\mathbb{R}_{>}/N^{2}$ is slower at special points $\Theta=n\pi$, which are also the cases where the two modes have the same rate with $\Gamma_{+}=\Gamma_{-}$. This originates mathematically from the fact that the overlap $c$ approaches 0 for $\Theta=n\pi$, which physically leads to two degenerate orthogonal modes.
@@
We note that, though the bound $\mathbb{R}_{<}$ in Sec. V.1 is a much tighter bound, it always has a finite-size correction $\sim\mathcal{O}(N^{-1})$. The reason could be that the product state ansatz is valid for large $N$, which therefore fails to capture the intermediate-size crossover.
Appendix B DTWA+TWA
From the Lindblad master equation (7), we can write the Heisenberg EOM for the spin and bosonic operators. Then, to obtain Eq. (9), we can treat the operators as complex numbers, using the discrete Wigner representation [72] for the spins and the Wigner representation [73] for the bosons. In this way, the quantum fluctuations of the initial state are incorporated into the EOMs, which can go beyond the MF regime.
B.1 Reduction of the SDE
$\begin{split}d\alpha^{(0)}_{\text{R}/\text{L}}&=-\frac{\kappa}{2}\alpha^{(0)}_{\text{R}/\text{L}}dt+\sqrt{\frac{\kappa}{2}}dW_{\text{R}/\text{L}},\end{split}$ (50)
and a deterministic part
[…]
For the initial state, we can let $\alpha^{(0)}_{\text{R}/\text{L}}(0)=\alpha_{\text{R}/\text{L}}(0)$ and $\alpha_{\text{R}/\text{L}}^{(1)}(0)=0$. The stochastic part can be solved analytically, or in practice with a tiny time step for later use. Then $\alpha^{(0)}_{\text{R}/\text{L}}$ enter the deterministic EOM of $\alpha^{(1)}_{\text{R}/\text{L}}$ in (51) and the deterministic EOM of spin variables in (9), which can then be solved to high precision with high-order integrators.
B.2 Adiabatic Elimination
In the Markovian limit, where the bosonic modes are heavily damped with $\kappa\gg\gamma N$, we can adiabatically eliminate the dynamics of $\alpha_{\text{R}}(t)$ and $\alpha_{\text{L}}(t)$ to derive a set of reduced stochastic equations for the spin variables only. With $\kappa/(\gamma N)\gg 1$, Eq. (9) can be formally integrated as
$\begin{split}\alpha_{\text{R}/\text{L}}(t)=\sqrt{\frac{\gamma}{2\kappa}}\tilde{J}_{\text{R}/\text{L}}(t)+\sqrt{\frac{2}{\kappa}}dW_{\text{R}/\text{L}}/dt.\end{split}$ (52)
This tells us that the occupation number in the cavity is of order $\mathcal{R}/\kappa$, which is $\sim\mathcal{O}(1)$ and $\sim\mathcal{O}(N)$ for $\mathcal{R}\sim\gamma N$ and $\mathcal{R}\sim\gamma N^{2}$ respectively when $\kappa\sim\gamma N$. Then, by putting (52) back to the EOM for the spins (9) and converting the obtained Stratonovich SDE to Ito SDE [74], we obtain
@@ B.2 Adiabatic Elimination
$\begin{split}ds_{j}^{-}=&-\frac{\gamma}{2}s_{j}^{-}+\frac{\gamma}{2}\sum_{l}s_{j}^{z}s_{l}^{-}e^{i|\xi_{j}-\xi_{l}|}+i\sqrt{\frac{\gamma}{2}}s_{j}^{z}dW_{j}\\ ds_{j}^{z}=&-\gamma s_{j}^{z}-\gamma\sum_{l}\left(s_{j}^{+}s_{l}^{-}e^{i|\xi_{j}-\xi_{l}|}+\text{c.c.}\right)\\ &-\sqrt{2\gamma}\left(s_{j}^{+}dW_{j}+\text{c.c.}\right),\end{split}$ (53)
@@
where the $dW_{j}=(e^{\xi_{j}}dW_{\text{R}}+e^{-\xi_{j}}dW_{\text{L}})/\sqrt{2}$ are a set of $N$ interdependent complex Wiener increments, defined in terms of the two independent increments $dW_{\text{R}}$ and $dW_{\text{L}}$. Note that these equations are valid on timescales that are large compared with $\kappa^{-1}$.
@@
From a numerical perspective, eliminating bosonic modes avoids a costly simulation of the bosonic amplitudes on a very fast timescale and enforces the Markovian limit assumed in the original master equation. Note that recently a similar derivation has been performed Ref. [58] using Keldysh formalism, which leads to the same stochastic EOMs as in Eq. (53). However, keeping the bosonic modes as dynamical degrees of freedom has the advantage that it can be readily extended to a non-Markovian setting, as discussed in Sec. VI.2.
@@ Appendix C Numerical Benchmarking of Different Methods
Here, in complement to Sec. III.4, we show numerical benchmarks of three approximate numerical methods – a variant of DTWA in spherical coordinates (sDTWA), quantum state diffusion with product state approximation (QSDMF), and cumulant expansion (CE) – against exact simulations using either quantum jump (QJ) or exact diagonalization (ED). In the Markovian case described by (1), as shown in Fig. 12 (a-c), all these methods are applicable and accurate for computing the superradiant peaks.
In the non-Markovian case, where $N=5,10,15$ atoms coupled with one lossy cavity of decay rate $\kappa=\gamma N$, only CE and our DTWA are applicable. We first benchmark our DTWA formalism in the study of $N$ homogeneous atoms coupled with one lossy cavity, where the permutation invariance permits exact simulations with the large spin basis for moderate system size. Since the cavity can be truncated to keep the first $(N+1)$ levels, the Hilbert space dimension needed for exact diagonalization (ED) is $(N+1)^{2}$. In this setting, we compute the collective decay rate $\mathcal{R}(t):=-\frac{d}{dt}\sum_{j}\braket{\sigma_{j}^{z}}/2$. As can be seen in Fig. 11(a), even when $\kappa=\gamma N$, results from DTWA agree with those from ED remarkably well. In contrast, CE with order $C=3$ still shows significantly large errors [see Fig. 12(d)]. We thus conclude that our DTWA formalism is the only one of these four methods that can deal with non-Markovian effects, and remains efficient and accurate.
Figure 12: Comparison of different methods. Error in computing $\mathcal{R}/(\gamma N^{2})$ using (a) sDTWA, (b) QSDMF, and (c) CE (with order $C=2$ or 33), benchmarked against QJ simulations, for system sizes $N=5,10,15$. Each is averaged over $10^{3}$ trajectories. (d) Comparison between ED and CE, for simulating $N=5,10,15$ atoms coupled with one lossy cavity with $\kappa=\gamma N$.

### v2

@@ VII Conclusion and Outlook
In summary, we investigated the fate of Dicke superradiance in atomic arrays subject to strong disorder. Using semiclassical methods tailored for collective emission, we numerically demonstrate that the characteristic $N^{2}$-scaling of the peak emission rate and the scaling $t_{\star}\sim\log(N)/N$ of the peak emission time remain remarkably robust, even when atomic positions are strongly disordered. In the extreme limit of a homogeneously distributed atomic ensemble, we found that the peak rate is only reduced by a constant factor of approximately $1/4$ and that there is no change in the asymptotic scaling behavior. However, disorder affects the finite-size scaling toward this limit, which explains the difficulty of exploring these effects with most of the previously employed numerical methods that are limited to small atom numbers.
@@
To gain physical understanding of our numerical findings on this robustness, we developed a variational product-state ansatz that yields a physically motivated upper bound on the collective decay rate. The bound $\mathbb{R}_{<}$ in Eq. (33) not only captures the observed superradiant scaling for disorder of varying strength, but also offers conceptual insight into the mechanism behind its robustness. Specifically, it reveals a form of disorder-induced self-organization: instead of synchronizing to align along a global direction as in the weakly disordered case, the atomic spins align coherently along locally determined directions set by the relative propagation phases. This form of spontaneous spin ordering always leads to constructive interference for collective emission, even under strong disorders. Using QSDMF simulations that can provide access to trajectory-level quantities, the spin ordering is confirmed numerically. Interestingly, this spin ordering also reveals the mechanism that underlies the asymmetric photon correlations reported in [28], since the two ordering patterns prefer emission to the left and the right, respectively.
@@
Looking forward, several promising directions emerge. While our current model involves two decay channels arising from the bidirectional waveguide, extensions to systems with multiple decay channels may reveal new forms of collective phenomena. Additionally, studying superradiance in engineered environments—such as squeezed reservoirs, structured bath, or chiral waveguides—may uncover richer phenomena of coherence and self-organization. The analytical and numerical frameworks introduced here are broadly adaptable for future studies of disorder, coherence, and collectivity in many-body quantum optics. Beyond its fundamental interest, the predicted robustness of these collective quantum effects against various forms of local disorders is also of significant practical importance for potential applications that rely on collective radiation effects. This includes the efficient generation of metrologically useful states [46, 47, 48, 49], the development of ultra-narrow superradiant lasers [43, 44, 45], as well as recent proposals for superradiant quantum batteries [74]. Our findings suggest that such applications exhibit significantly greater resilience to imperfections than commonly assumed.
@@ Acknowledgements.
Appendix A Finite-Size Scaling of the Decay Rates
The total collective decay rate $\mathcal{R}$ introduced in Eq. (III.3), can in general be written as
@@ Appendix A Finite-Size Scaling of the Decay Rates
where $\hat{R}$ is a Hermitian matrix defined by the jump operators in Eq. (3), $\Gamma_{k}$ are its eigenvalues, and $\Omega_{k}=\sum_{j}U_{kj}\sigma_{j}^{-}$ are the corresponding decay modes after diagonalization of $\hat{R}$ with a unitary transformation $U$. While the eigenvalues are only directly related to the decay rates of the first few emissions, an upper bound of the maximal decay rate can be written out as (see Appendix E)
$\mathcal{R}_{\star}\leq\frac{3}{2}N\Gamma_{+}.$ (42)
Therefore, the upper bound indicates that $\mathcal{R}_{\star}\sim N\Gamma_{+}$, which means the finite-size scaling of $\mathcal{R}_{\star}$ could be hinted in the finite-size scaling of $\Gamma_{+}$. Though it is a priori unclear if $\mathcal{R}_{\star}\sim N\Gamma_{+}$ holds, our numerical results confirm so.
@@
of which the diagonalization gives two eigenvalues $\Gamma_{\pm}=\frac{\gamma}{2}N(1\pm|c|)$. Then our original master equation (1) can be rewritten as Eq. (5) with two orthogonal modes.
Since $\xi_{j}$ is uniformly distributed in $\Theta/2,\Theta/2)$, we know $\Re e^{-i2\xi}$ and $\Im e^{-i2\xi}$ have the means and variances shown in Eq. (19). Thus, for large $N$, $c$ obeys a Gaussian distribution with $\Re c\sim\mathcal{N}(\mu_{a},\sigma_{a}^{2}/N)$ and $\Im c\sim\mathcal{N}(\mu_{b},\sigma_{b}^{2}/N)$. In the thermodynamic limit $N\rightarrow\infty$, we then obtain Eq. (20).
To evaluate the finite-size scaling, we consider large but finite $N$. To get an analytical understanding, we can write $\Re c=\mu_{a}+\delta_{a}$ and $\Im c=\delta_{b}$, with $\delta_{a/b}$ being small Gaussian fluctuations $\sim N^{-1/2}$. Then we see that
@@
For the values of $\Theta=n\pi$, where $\mu_{a}=0$, we find $|c|=\sqrt{(\delta_{a}^{2}+\delta_{b}^{2})}\sim|\delta_{a/b}|\sim N^{-1/2}$. This explains the scaling $E[\Gamma_{+}]=\frac{1}{2}\gamma N\left(1+\mathcal{O}\left(N^{-1/2}\right)\right)$ for the finite-size corrections, in accordance with the exact results in Eq. (21). For any other $\Theta\neq n\pi$, where $\mu_{a}\neq 0$, we can make the expansion
$|c|=\mu_{a}\left(1+\frac{\delta_{a}}{\mu_{a}}+\frac{1}{2}\frac{\delta_{b}^{2}}{\mu_{a}^{2}}+\ldots\right)$ (47)
to obtain
$E[|c|]=\mu_{a}+\frac{1}{2\mu_{a}}E[\delta_{b}^{2}]\sim\mu_{a}+\mathcal{O}(N^{-1}).$ (48)
Therefore, in this case we find a finite-size correction of the form
$E[\Gamma_{+}]=\frac{1}{2}\gamma N\left(1+\mu_{a}+\mathcal{O}\left(N^{-1}\right)\right),$ (49)
consistent with Eq. (20).
@@
Therefore, the convergence is slower at special points $\Theta=n\pi$, where $\Gamma_{+}=\Gamma_{-}$. This originates mathematically from the fact that the overlap $c$ approaches 0 for $\Theta=n\pi$. We note that, though the bound $\mathbb{R}_{<}$ in Sec. V.1 is much tighter, it always has a finite-size correction $\sim\mathcal{O}(N^{-1})$. The reason could be that the product state ansatz is valid for large $N$, which therefore fails to capture the intermediate-size crossover. A similar difference in the finite-size scaling is numerically observed for the actual peak decay rate $\mathcal{R}_{\star}$, indicating the validity of the relation $\mathcal{R}_{\star}\sim N\Gamma_{+}$.
Appendix B Markovian Limit of DTWA
In Sec. III.1, we have presented the stochastic EOM for a set of TLAs coupled to two damped cavity modes. In numerical simulations, we can then choose a sufficiently large damping rate $\kappa$ to recover an effective Markovian dynamics for the TLAs only. When only Markovian dynamics is concerned, an adiabatic elimination of the cavity modes can be carried out analytically, as discussed in the following. From a numerical perspective, eliminating the cavity modes avoids a costly simulation of the bosonic amplitudes on a very fast timescale and enforces the Markovian limit assumed in the original master equation. However, keeping the bosonic modes as dynamical degrees of freedom has the advantage that it can be readily extended to a non-Markovian setting, as discussed in Sec. VI.2.
We start by introducing a time increment $\Delta t$, which is short with respect to the dynamics of the TLAs, but long compared to $\kappa^{-1}$. By integrating the EOM for the spin variables in Eq. (9) over this time interval, we obtain, for example,
@@ Appendix B Markovian Limit of DTWA
$\begin{split}\Delta s_{j}^{-}(t)&\simeq i\frac{\gamma}{2}\sum_{l\neq j}\sin(|\xi_{j}-\xi_{l}|)s_{j}^{z}s_{l}^{-}\Delta t\\ &+\frac{g}{\sqrt{2}}s_{j}^{z}\left(\Delta\alpha_{\text{R}}(t)e^{i\xi_{j}}+\Delta\alpha_{\text{L}}(t)e^{-i\xi_{j}}\right),\end{split}$ (50)
where $\Delta\alpha_{\text{R,L}}(t)=\int_{t}^{t+\Delta t}\alpha_{\text{R,L}}(s)\,ds$ are the corresponding increments of the cavity amplitudes. From the formal solutions of Eq. (10), we have
$\begin{split}\alpha_{\text{R,L}}(t)=&\frac{g}{\sqrt{2}}\int_{0}^{t}ds\,e^{-\frac{\kappa}{2}(t-s)}\tilde{J}_{\rm R,L}(s)\\ &+\sqrt{\frac{\kappa}{2}}\int_{0}^{t}e^{-\frac{\kappa}{2}(t-s)}dW_{\rm R,L}(s)\\ \approx&\frac{\sqrt{2}g}{\kappa}\tilde{J}_{\rm R,L}(t)+\sqrt{\frac{2}{\kappa}}\Xi_{\rm R,L}(t),\end{split}$ (51)
where the last approximation is valid when $\kappa/(\gamma N)\gg 1$. Here, we have introduced the noise process
$\Xi_{\rm R,L}(t)=\frac{\kappa}{2}\int_{0}^{t}e^{-\frac{\kappa}{2}(t-s)}dW_{\rm R,L}(s),$ (52)
which is $\delta$-correlated on timescales longer than $\kappa^{-1}$. From Eq. (51), we see that the occupation number in the cavity is of order $\mathcal{R}/\kappa$, which is $\sim\mathcal{O}(1)$ and $\sim\mathcal{O}(N)$ for $\mathcal{R}\sim\gamma N$ initially and $\mathcal{R}\sim\gamma N^{2}$ around superradiance respectively, when $\kappa\sim\gamma N$. By defining the corresponding coarse-grained increments, $\Delta W_{\rm R,L}(t)=\int_{t}^{t+\Delta t}\Xi_{\rm R,L}(s)$, we obtain the effective stochastic updates for the spin variables,
$\begin{split}\Delta s_{j}^{-}(t)\simeq&\frac{\gamma}{2}\sum_{l}s_{j}^{z}s_{l}^{-}e^{i|\xi_{j}-\xi_{l}|}\Delta t\\ &+\sqrt{\frac{\gamma}{4}}s_{j}^{z}\left(\Delta W_{\rm R}(t)e^{i\xi_{j}}+\Delta W_{\rm L}(t)e^{-i\xi_{j}}\right),\end{split}$ (53)
together with a similar equation for the $\Delta s_{j}^{z}(t)$. On the timescale of the spin variables, we can now treat all increments as infinitesimal to recover a regular set of stochastic differential equations. These are, however, of Stratonovich type and must be converted into Ito form [76] for numerical simulations. We finally end up with
$\begin{split}ds_{j}^{-}=&-\frac{\gamma}{2}s_{j}^{-}+\frac{\gamma}{2}\sum_{l}s_{j}^{z}s_{l}^{-}e^{i|\xi_{j}-\xi_{l}|}+\sqrt{\frac{\gamma}{2}}s_{j}^{z}dW_{j}\\ ds_{j}^{z}=&-\gamma s_{j}^{z}-\gamma\sum_{l}\left(s_{j}^{+}s_{l}^{-}e^{i|\xi_{j}-\xi_{l}|}+\text{c.c.}\right)\\ &-\sqrt{2\gamma}\left(s_{j}^{+}dW_{j}+\text{c.c.}\right),\end{split}$ (54)
@@
where the $dW_{j}=(e^{\xi_{j}}dW_{\text{R}}+e^{-\xi_{j}}dW_{\text{L}})/\sqrt{2}$ are a set of $N$ interdependent complex Wiener increments, defined in terms of the two independent increments $dW_{\text{R}}$ and $dW_{\text{L}}$.
@@ Appendix C Numerical Benchmarking of Different Methods
Figure 12: Comparison of different numerical methods. Error in computing $\mathcal{R}/(\gamma N^{2})$ using (a) sDTWA, (b) QSDMF, and (c) CE (with order $C=2$ or 33), benchmarked against QJ simulations, for system sizes $N=5,10,15$. All averages have been taken over $\mathcal{N}=10^{3}$ trajectories. (d) Comparison between ED and CE for simulating $N=5,10,15$ atoms coupled with one lossy cavity with $\kappa=\gamma N$.
Here, in complement to Sec. III.4, we show numerical benchmarks of three approximate numerical methods – a variant of DTWA in spherical coordinates (sDTWA), quantum state diffusion with product state approximation (QSDMF), and cumulant expansion (CE) – against exact simulations using either quantum jump (QJ) or exact diagonalization (ED). In the Markovian case described by Eq. (1), as shown in Fig. 12 (a-c), all these methods are applicable and accurate for computing the superradiant peaks.
For the non-Markovian case, we consider $N$ TLA that are coupled uniformly to one lossy cavity with decay rate $\kappa=\gamma N$, such that the permutation symmetry permits exact simulations with moderate atom numbers of $N=5,10,15$. Since the cavity can be truncated to keep the first $(N+1)$ levels, the Hilbert space dimension needed for exact diagonalization (ED) is $(N+1)^{2}$. In this setting, we compute the collective decay rate as $\mathcal{R}(t)=-\frac{d}{dt}\sum_{j}\braket{\sigma_{j}^{z}}/2$. As can be seen in Fig. 11(a), even when $\kappa=\gamma N$, results from DTWA agree with those from ED remarkably well. In contrast, CE with order $C=3$ still shows significantly larger errors [see Fig. 12(d)]. We thus conclude that our DTWA formalism is the only one of these four methods that can deal with non-Markovian effects, while remaining efficient and accurate.